OpenStudy (anonymous):

A designer wants to make use of a half-spherical cistern in a water system. If the cistern has a diameter of 6.5 feet, how many gallons of water can it hold?

OpenStudy (campbell_st):

well the volume of a sphere is \[V = \frac{4}{3} \pi r^3\] you know r = 6.5 ft so the volume of a half sphere is found by dividing the volume calculated above by 2. then convert cubic feet to capacity... gallons... 1ft^3 = 7.48052 gal

OpenStudy (goformit100):

"Welcome to OpenStudy. I can guide regarding this useful site; ask your doubts from me, for it you can message me. Please use the chat for off topic questions. And remember to give the helper a medal, by clicking on "Best Answer". We follow a code of conduct, ( http://openstudy.com/code-of-conduct ). Please take a moment to read it."